• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 07:05:48 on February 1,2008

    Entry number 228532

    keyword=CSR test: Owl shift summary

    Owl shift for CSR Pb taregt cross calibration:
    Shift workers: J. P. Chen (SL), X. Yan (TO), Y. Oh (3rd)
    MCC vacuum problem earlier, PREX test requested to extend 
    23:30 started to move right spectrometer to 25 degree and set right magnets to 
    958.4 MeV/c. PREX lumi test continues.
    midnight: spectrometer set. Requested raster size chaneg to 1.5x 4 mm.
    And position to (-3.5,0). MCC were able to get BPMB to -3.5 mm, but
    BPMA only to -2.2 mm.
    0:00 started data taking with thin Pb target. Left DAQ problem: only can take 
    ~300 Hz. Alexandre is checking, but problem persists.
    Moved to Thick PREX Pb target.
    Moved to loop 3 Pb+LH2 target.
    1:15 page/called Bob Michaels.
    Bob gave phone instruction for testing out.
    1:45 changed to 30 degrees for both spectrometers.
    2:00 Bob came in to work on fixing the problem with Left DAQ.
    2:35 Bob fixed left DAQ problem
    2:50 Alexandre lowered HV (~400 V) for the last 18 channels (LeCroy)
    for NaI detector 
    3:00 Alex/Bob rebooted adaql1 to get halog work, which affected
    DAQ (left stuck 6075 and right got high deadtime 3747)
    4:15 move to thin Pb.
    Right DAQ rate drops after a while. Not sure why. Stop run and start
    again, it becomes fine.
    5:17 moved to thick Pb.
    6:00 changed to 25 degree and P to 982.36 MeV (+2.5%) for both arms.
    moved to loop 3 Pb.
    6:08 right dipole set.
    6:25 move to thin Pb.
    6:25 started to slowly closing down J-T for loop 3 and tuning down fan speed.
    6:30-6:35 MCC doing QE measurement.
    6:45 move to thick Pb.
    7:00 stop beam. End of CSR test.
    requested controlled access and started warming up target.
    P=958.4 MeV/c
    Angle    Target		Right     	Left
    25	 thin Pb	3741/3742	6054 (DAQ problem)
    25	 thick Pb       3743/3744	6055-6060? (DAQ problem)
    25	 Loop 3 Pb	3745		6062-6064 (DAQ problem)
    30       loop 3 Pb	3746/3747*	6065-6075 (DAQ problem)	
    			3748 (good run)	6076/6077 (good run)
    30       thin Pb		3749/3750	6078/6079	
    30       thick Pb	3751/3752	6080
    changed P to 982.36 MeV (+2.5%)
    25       Loop 3 Pb	3753 (dipole not set) 6081
    			3754 (dipole set)
    25	thin Pb		3755		6082
    25	thick Pb	3756		6083/6084